The frequency of a fundamental note from a closed pipe is 200Hz. What is the frequency of the next possible note from the same pipe?
1
Expert's answer
2020-09-10T14:40:11-0400

The fundamental frequency (with "n = 1" ) in a closed pipe is given by the following expression:


"f_1 = \dfrac{nv}{2L} = \dfrac{v}{2L}"

where "v = 340 m\/s" is the speed of sound in air. Then the length of the pipe will be:


"L = \dfrac{v}{2f_1}"

The next possible note will appear at "n = 2" .Then:


"f_2 = \dfrac{2v}{2L} = 2f_1 = 2\cdot 200 Hz = 400Hz"

Answer. 400 Hz.